# The topology of a chaotic attractor in the Kuramoto-Sivashinsky equation

### v1

@@ Abstract
The Birman-Williams theorem gives a connection between the collection of unstable periodic orbits (UPOs) contained within a chaotic attractor and the topology of that attractor, for three-dimensional systems. In certain cases, the fractal dimension of a chaotic attractor in a partial differential equation (PDE) is less than three, even though that attractor is embedded within an infinite-dimensional space. Here we study the Kuramoto-Sivashinsky PDE at the onset of chaos. We use two different dimensionality-reduction techniques – proper orthogonal decomposition and an autoencoder neural network – to find two different approximate embeddings of the chaotic attractor into three dimensions. By finding the projection of the attractor’s UPOs in these reduced spaces and examining their linking numbers, we construct templates for the branched manifold which encodes the topological properties of the attractor. The templates obtained using two different dimensionality reduction methods mirror each other. Hence, the organization of the periodic orbits is identical (up to a global change of sign) and consistent symbolic names for low-period UPOs are derived. This is strong evidence that the dimensional reduction is robust, in this case, and that an accurate topological characterization of the chaotic attractor of the chaotic PDE has been achieved.
@@ I Introduction
The importance of unstable periodic orbits (UPOs) to the understanding of chaotic dynamics has long been appreciated. It is believed, and in some cases proven, that UPOs are dense within strange attractors, and so by finding all the UPOs and a suitable weighting, it should be possible to average their properties to find statistics for the chaotic dynamics. This is the essence of ‘periodic orbit theory’, which has been successfully applied in simple systems (Artuso, Aurell, and Cvitanovic, 1990; Lan, 2010), and in particular in three-dimensional systems of coupled ordinary differential equations (ODEs) (Eckhardt and Ott, 1994; Maiocchi, Lucarini, and Gritsun, 2022). Increasingly, effort is being made to apply the ideas of periodic orbit theory to infinite-dimensional partial differential equation (PDE) systems, and in particular to the problem of turbulence in fluid dynamics (Chandler and Kerswell, 2013; bud, 2017; McCormack, Cavalieri, and Hwang, 2024). However, the collection of periodic orbits can tell us much more about the dynamics than only its statistics, such as the configuration of the attractor in state space.
@@
One approach to understand the behaviour of chaotic attractors is through ideas from topology. In three dimensions, UPOs, which are closed loops in state space, form knots, and pairs of UPOs form links. Such knots and links possess topological invariants, which can be used to analyse the topological structure of the underlying chaotic attractor. In particular, the Birman-Williams theorem (Birman and Williams, 1983a, b) shows that a chaotic attractor in a three-dimensional system can be identified with a two-dimensional branched manifold, a so-called ‘template’, on which all periodic orbits lie. The structure of this branched manifold – the number of branches and how they link together – can be deduced from the linking numbers of the periodic orbits. This formalises the observation that chaotic attractors in canonical examples such as the Lorenz and Rössler systems seem to be very thin, almost two-dimensional structures, and the templates found for these systems match the structures visible in a long chaotic trajectory in the usual coordinates. The method of determining templates for all periodic orbits has been successfully applied to many three-dimensional ODE systems, including ones for which the topology is not obvious from observing trajectories (Birman and Williams, 1983a; Letellier, Dutertre, and Maheu, 1995; Letellier and Aguirre, 2012).
@@
In four and higher dimensions, all UPOs are however homeomorphic to the unknot, and thus the specific methods discussed here cannot be directly applied. The solution spaces of PDEs are infinite dimensional. In practice, when solving PDE dynamical systems numerically, this infinite dimensional space is discretised to be finite-dimensional, but the number of dimensions is still on the order of at least tens, if not much higher, and so the ideas of Birman and Williams are not directly applicable (although more modern approaches exist which can be applied to dimensions higher than three, as discussed in section V). Despite the high-dimensional solution spaces, the long-time dynamics of many systems often collapses onto an attracting set of much lower dimension.
@@
The dissipative PDE systems that arise in fluid dynamics are typically associated with a parameter, such as the Rayleigh number or Reynolds number, whose increase is correlated with an increase in the complexity of the dynamics. When such a parameter is sufficiently high the system is fully turbulent, as characterised by a large number of positive Lyapunov exponents. In this case, the global attractor is a relatively high- (but still finite-) dimensional fractal structure in state space. In contrast, when the parameter is sufficiently small, the system collapses onto a single attracting fixed point. Between these regimes, it is often possible to find chaotic dynamics with only one positive Lyapunov exponent, which indicates a strange attractor with less than three dimensions (in the absence of symmetries), embedded within the infinite-dimensional state space of the PDE. If it were possible to construct an embedding of this low dimensional attractor within a three-dimensional space, this would preserve the topological properties, and we would be able to use the ideas of Birman and Williams to find a template for the system.
@@
Since the topological properties of dynamical systems are independent of the choice of coordinates used for parameterising state space, they provide a way to compare different representations of the same or similar systems. This is particularly relevant to PDE systems, in which the dynamics are necessarily discretised in order to be solved on a computer. Different choices of discretisation (for example, finite differences or Galerkin methods) may lead to very different state space geometries, but the topology must be equivalent if the discretisations are accurate. Furthermore, it is often desirable to find a ‘reduced-order-model’, a small set of ordinary differential equations approximating the dynamics of a PDE system, either to increase computational efficiency or to aid the interpretation of physical mechanisms underlying the dynamics. One way to validate the accuracy of such a model would be to check that the model preserves the topology of the dynamics of the full system. However, in order for this we need a method to distill the topology of a chaotic PDE, which does not currently exist.
@@
The idea that autoencoders can preserve the topology of a chaotic attractor is supported by the studies of Uribarri and Mindlin (2020) and Fainstein et al. (2023), who applied autoencoders to time-series from synthetic and experimental data. Unlike in the present work, the aim there was to determine whether the known topological properties are preserved so that the autoencoders give an embedding. Indeed, Fainstein et al. (2023) showed that this is not always the case, and so autoencoders should be used with caution.
@@ II The Kuramoto-Sivashinsky equation
Equation 1 is invariant under the transformation $x\mapsto-x$, $u\mapsto-u$ and therefore admits an invariant subspace of solutions for which $u(-x)=-u(x)$. Following previous studies (Cvitanović, Davidchack, and Siminos, 2010; Lasagna, 2018), we consider the dynamics constrained to this (usually unstable) subspace. This has the advantage of removing the continuous symmetry $x\mapsto x+\Delta$, which otherwise would increase the dimension of any $x$-varying attractor by 1, significantly complicating topological analysis. In this subspace, the first known chaotic attractor occurs at $L\approx 18.05$. Wilczak and Zgliczyński (2020) give a computer assisted proof for the existence of this attractor at $L=2\pi/\sqrt{0.1212}=18.048\dots$, which is the parameter value we use. Indeed, they proved the existence of a symbolic dynamics for the system, which in principle allows the enumeration and computation of a countable infinity of UPOs. We found, empirically, that this attractor has a very small basin of attraction: it is necessary to use an initial condition very close to the states depicted in Wilczak and Zgliczyński (2020), otherwise the system rapidly collapses onto a fixed point.
@@
In order to be able to embed the chaotic attractor in a space of dimension 3, it is necessary but not sufficient for the fractal dimension to be less than three. Consider for example the Klein bottle, whose dimension is only two but cannot be embedded within a Euclidean space of dimension less than four. Experiments at larger values of $L$, for example the well-studied(Lasagna, 2018) $L\approx 39$, and the parameter values studied by Siminos (2021) of $L\approx 36$, found attractors whose fractal dimension was apparently less than three, but for which we could only produce a non-injective immersion into a 3D space. An injective embedding is necessary for our topological approach to be valid.
@@ II.1 Numerical solution
We discretize the spatial dimension of the anti-symmetric system with $N_{x}=32$ points, turning the function $u$ into a 32-dimensional vector $\bm{u}$. We observe a satisfactory drop in the amplitudes of the Fourier modes for this discretization, suggesting a well-resolved discrtization of the PDE. For comparison, Lan and Cvitanović (2008) also use 32 modes in the $L=38.5$ system to find periodic orbits, which exhibits stronger chaos (Edson et al., 2019) and smaller-scale spatial features than our system. To calculate the POD modes in section III.1 and train the autoencoder in section III.2, we generate a long direct numerical simulation (DNS) of the system using the ETDRK4 scheme(Kassam and Trefethen, 2005), with $t_{\mathrm{max}}=20000,dt=0.1$. Due to the small basin of attraction of the chaotic attractor, we pick a point that we know for certain lies on the attractor as initial condition for the DNS.
A section of this timeseries is shown in figure 1, in the typical manner of plotting space along the vertical axis and using a colour scale to show the value of $u$ as time varies. Unlike usual projection of plotting trajectories in the Lorenz system, for example, no topology for the dynamics is evident in this figure. Indeed, for this value of $L$ it is not even obvious that the dynamics is chaotic.
@@ III.1 Proper orthogonal decomposition
The most common dimensionality-reduction techniques are linear methods based on eigen-analysis such as dynamic mode decomposition (DMD) (Schmid, 2022) and principal component analysis (PCA) (Pearson, 1901; Hotelling, 1936), better known as proper orthogonal decomposition (POD) in the fluid dynamics community. POD consists of determining the dominant (uncorrelated) modes that capture most of the system’s variance, or ‘kinetic energy’. Given a time-series $\{\bm{u}_{i}\}_{i=1}^{r}$ with $r$ time-steps, where $\bm{u}_{i}\in\mathbb{R}^{N_{x}}$ we can calculate the POD modes by stacking the zero-mean time-series $\bm{\tilde{u}}_{i}=\bm{u}_{i}-\bm{u_{\mathrm{mean}}}$ in a matrix $\bm{\tilde{U}}\in\mathbb{R}^{r\times N_{x}}$ (where the rows are $r$ time-steps), and by considering its covariance matrix. The unbiased estimator $\bm{C}\in\mathbb{R}^{N_{x}\times N_{x}}$ for the covariance matrix is given by
@@ III.2 Autoencoder
Since we define the latent dimension to be $N_{h}=3$, this gives us the desired 3-dimensional representation of the KSE, in which we would like to test the methods described in section IV.1. Taking datapoints from a long physical time-series and applying the encoder, we obtain a 3D latent time-series. Figure 6 shows a 3D plot of the image of the attractor in latent space, illustrating its band-like shape, and 2D projections.
Figure 6: 3D plot (top) and 2D projections (bottom) of the chaotic attractor in the autoencoder’s latent space. The latent coordinates $(h_{1},h_{2},h_{3})$ are obtained by applying the encoder $\mathcal{E}$ to a data point. The grey points are from a time-series solving the PDE. The blue line is a periodic orbit with period $T\approx 51.62$, plotted to emphasise the shape of the attractor.
@@ III.3.1 POD return map
In the POD latent coordinates $(\xi_{1},\xi_{2},\xi_{3})$, we define a Poincaré section $\mathcal{P}_{\mathrm{POD}}$ given by the plane $\xi_{1}=1.3$ and flow direction $\partial_{t}\xi_{1}<0$. The value $\xi_{1}=1.3$ is chosen so that the flow is approximately orthogonal to the Poincaré section. This defines a unique crossing section along the band (see figure 7), and gives us a sequence $\{\xi_{2}^{(i)}\}$ of the $\xi_{2}$ coordinate of the crossings through $\mathcal{P}_{\mathrm{POD}}$. The return map $f_{1}$ is defined by $f_{1}:\xi_{2}^{(i)}\mapsto\xi_{2}^{(i+1)}$. A plot of such subsequent crossings is shown in the top left panel of figure 8. The points approximately fall on a smooth unimodal curve, which we fit with a polynomial of order three. This simplified first return map appears to only have one intersection with the identity $\xi_{2}^{(i)}=\xi_{2}^{(i+1)}$, suggesting that there is only one periodic orbit with $p=1$, where $p$ is the number of intersections of an orbit with $\mathcal{P}_{\mathrm{POD}}$. By looking at further iterations/compositions of the return map (see other panels of figure 8), we also obtain an indication that there are 1, 0 and 1 periodic orbits with $p=2,3,4$ respectively (accounting for crossings that correspond to periodic orbits with smaller $p$ e.g. the orbit with $p=1$ will also account for one intersection in all subsequent graphs). Finally, we split the band domain $I=[-3.229,-2.952]$ at the minimum value of the first return map, giving approximately $I_{0}=[-3.229,-3.055]$ and $I_{1}=[-3.055,-2.952]$. From this, we obtain an indication of the qualitative dynamics: points starting in $I_{0}$ are mapped to the full interval $I$, while points starting in $I_{1}$ are mapped to a subset of $I_{0}$. This is shown in fig. 9.
Figure 7: $(\xi_{1},\xi_{2})$ projection of the chaotic attractor in the latent space defined by the POD coordinates. The grey points are from a physical time-series that is put through the encoder. The magenta line illustrates the Poincaré section $\xi_{1}=1.3$ and $\partial_{t}\xi_{1}<0$, denoted $\mathcal{P}_{\mathrm{POD}}$.
@@
Figure 9: First return map for the Poincaré section $\mathcal{P}_{\mathrm{POD}}$. The blue dots show data from a chaotic timeseries, whereas the green dot, red line and black line show the periodic orbits $1_{1}$, $2_{1}$ and $4_{1}$ respectively (see table 1). Below the figure, the intervals represent the first and second level within the state space partition derived from the simplified return map shown in figure fig. 11. These intervals are used to construct symbolic names for the periodic orbits.

### v2

@@ Abstract
The Birman-Williams theorem gives a connection between the collection of unstable periodic orbits (UPOs) contained within a chaotic attractor and the topology of that attractor, for three-dimensional systems. In certain cases, the fractal dimension of a chaotic attractor in a partial differential equation (PDE) is less than three, even though that attractor is embedded within an infinite-dimensional space. Here we study the Kuramoto-Sivashinsky PDE at the onset of chaos. We use two different dimensionality-reduction techniques – proper orthogonal decomposition and an autoencoder neural network – to find two different mappings of the chaotic attractor into three dimensions. By finding the image of the attractor’s UPOs in these reduced spaces and examining their linking numbers, we construct templates for the branched manifold which encodes the topological properties of the attractor. The templates obtained using two different dimensionality reduction methods are equivalent. The organization of the periodic orbits is identical and consistent symbolic sequences for low-period UPOs are derived. While this is not a formal mathematical proof, this agreement is strong evidence that the dimensional reduction is robust, in this case, and that an accurate topological characterization of the chaotic attractor of the chaotic PDE has been achieved.
@@ I Introduction
The importance of unstable periodic orbits (UPOs) to the understanding of chaotic dynamics has long been appreciated. It is believed, and in some cases proven, that UPOs are dense within strange attractors. Modern computational methods enable the detection and convergence of large numbers of such UPOs in high-dimensional discretizations of dynamical systems governed by partial differential equations (PDEs), including the Navier-Stokes equations from fluid dynamics (Budanur et al., 2017; Parker and Schneider, 2022; Page et al., 2024; McCormack, Cavalieri, and Hwang, 2024). Especially in the context of fluid turbulence, UPOs carry the hope that we may be able to describe statistical properties of the turbulent flow as weighted ensemble averages over these UPOs. This envisioned form of ‘periodic orbit theory’ would transfer well-studied ergodic theory from low-dimensional chaotic ordinary differential equation (ODE) systems (Cvitanović et al., 2016; Lan, 2010; Eckhardt and Ott, 1994) to PDEs.
@@
One alternative approach to use periodic orbits for understanding the behaviour of chaotic attractors is through ideas from topology. In three dimensions, UPOs, which are closed loops in state space, form knots, and pairs of UPOs form links. Such knots and links possess topological invariants, which can be used to analyse the topological structure of the underlying chaotic attractor. In particular, the Birman-Williams theorem (Birman and Williams, 1983a, b) shows that a chaotic attractor in a three-dimensional system can be identified with a two-dimensional branched manifold, a so-called ‘template’, on which all periodic orbits lie. The structure of this branched manifold – the number of branches and how they link together – can be deduced from the linking numbers of the periodic orbits. This formalises the observation that chaotic attractors in canonical examples such as the Lorenz and Rössler systems seem to be very thin, almost two-dimensional structures, and the templates found for these systems match the structures visible in a long chaotic trajectory in the usual coordinates. The method of determining templates for all periodic orbits has been successfully applied to many three-dimensional ODE systems, including ones for which the topology is not obvious from observing trajectories (Birman and Williams, 1983a; Letellier, Dutertre, and Maheu, 1995; Letellier and Aguirre, 2012).
@@
In four and higher dimensions, all UPOs are however equivalent (i.e. ambient isotopic) to the unknot, and thus the specific methods discussed here cannot be directly applied. The solution spaces of PDEs are infinite dimensional. In practice, when solving PDE dynamical systems numerically, this infinite dimensional space is discretized to be finite-dimensional, but the number of dimensions is still on the order of at least tens, if not much higher, and so the ideas of Birman and Williams are not directly applicable (although more modern approaches exist which can be applied to dimensions higher than three, as discussed in section V). Despite the high-dimensional solution spaces, the long-time dynamics of many systems often collapses onto an attracting set of much lower dimension.
@@
The dissipative PDE systems that arise in fluid dynamics are typically associated with a parameter, such as the Rayleigh number or Reynolds number, whose increase is correlated with an increase in the complexity of the dynamics. When such a parameter is sufficiently high the system is fully turbulent, as characterised by a large number of positive Lyapunov exponents. In this case, the global attractor is a relatively high- (but still finite-) dimensional fractal structure in state space. In contrast, when the parameter is sufficiently small, the system collapses onto a single attracting fixed point. Between these regimes, it is often possible to find chaotic dynamics with only one positive Lyapunov exponent, which indicates a strange attractor with less than three dimensions (in the absence of symmetries), embedded within the infinite-dimensional state space of the PDE. If it were possible to construct an embedding of this low dimensional attractor into three-dimensional space, this would preserve the topological properties, and we would be able to use the ideas of Birman and Williams to find a template for the system.
@@
Since the topological properties of dynamical systems are independent of the choice of coordinates used for parameterising state space, they provide a way to compare different representations of the same or similar systems. This is particularly relevant to PDE systems, in which the dynamics are necessarily discretized in order to be solved on a computer. Different choices of discretization (for example, finite differences or Galerkin methods) may lead to very different state space geometries, but the topology must be equivalent if the discretizations are accurate. Furthermore, it is often desirable to find a ‘reduced-order-model’, a small set of ordinary differential equations approximating the dynamics of a PDE system, either to increase computational efficiency or to aid the interpretation of physical mechanisms underlying the dynamics. One way to validate the accuracy of such a model would be to check that the model preserves the topology of the dynamics of the full system. However, in order for this we need a method to distill the topology of a chaotic PDE, which does not currently exist.
Previous authors have attempted to find reduced-order models with the same dynamics as infinite dimensional systems, for example by using global modeling to model the dynamics with a three-dimensional ODE, from which a template can be derived (Ghosh et al., 2017). Our approach is instead to consider accurate periodic trajectories of the full PDE system, and then to embed these trajectories in a three-dimensional reduced space.
@@
The idea that autoencoders can in some cases preserve the topology of a chaotic attractor is supported by the studies of Uribarri and Mindlin (2020) and Fainstein et al. (2023), who applied autoencoders to timeseries from synthetic and experimental data. Unlike in the present work, the aim there was to determine whether the known topological properties are preserved so that the autoencoders give an embedding. Indeed, both of these papers showed that this is not always the case, and so autoencoders should be used with caution.
@@ II The Kuramoto-Sivashinsky equation
Equation 1 is invariant under the transformation $x\mapsto-x$, $u\mapsto-u$ and therefore admits an invariant subspace of solutions for which $u(-x)=-u(x)$. Following previous studies(Cvitanović, Davidchack, and Siminos, 2010; Lasagna, 2018; Dong, 2018), we consider the dynamics constrained to this (usually unstable) subspace. This has the advantage of removing the continuous symmetry $x\mapsto x+\Delta$, which otherwise would increase the dimension of any $x$-varying attractor by 1, significantly complicating topological analysis. In this subspace, the first known chaotic attractor occurs at $L\approx 18.05$. Wilczak and Zgliczyński (2020) give a computer assisted proof for the existence of this attractor at $L=2\pi/\sqrt{0.1212}=18.048\dots$, which is the parameter value we use. Indeed, they proved the existence of a symbolic dynamics for the system, which in principle allows the enumeration and computation of a countable infinity of UPOs. We found, empirically, that this attractor has a very small basin of attraction: it is necessary to use an initial condition very close to the states depicted in Wilczak and Zgliczyński (2020), otherwise the system rapidly collapses onto a fixed point.
@@
In order to be able to embed the chaotic attractor in a space of dimension 3, it is necessary but not sufficient for the fractal dimension to be less than three. Consider for example the Klein bottle, whose dimension is only two but cannot be embedded within a Euclidean space of dimension less than four. Experiments at larger values of $L$, for example the well-studied(Lasagna, 2018) $L\approx 39$, and the parameter values studied by Dong (2018) and Siminos (2021) of $L\approx 36$, found attractors whose fractal dimension was apparently less than three, but for which we could only produce a non-injective immersion into a 3D space. An injective embedding is necessary for our topological approach to be valid.
@@ II.1 Numerical solution
We discretize the spatial dimension of the anti-symmetric system with $N_{x}=32$ points, turning the function $u$ into a 32-dimensional vector $\bm{u}$. We observe a satisfactory drop in the amplitudes of the Fourier modes for this discretization, suggesting a well-resolved discretization of the PDE. For comparison, Lan and Cvitanović (2008) also use 32 modes in the $L=38.5$ system to find periodic orbits, which exhibits stronger chaos (Edson et al., 2019) and smaller-scale spatial features than our system. To calculate the POD modes in section III.1 and train the autoencoder in section III.2, we generate a long direct numerical simulation (DNS) of the system using the ETDRK4 scheme(Kassam and Trefethen, 2005), with $t_{\mathrm{max}}=20000,dt=0.1$. Due to the small basin of attraction of the chaotic attractor, we pick a point that we know for certain lies on the attractor as initial condition for the DNS.
A section of this timeseries is shown in figure 1, in the typical manner of plotting space along the vertical axis and using a colour scale to show the value of $u$ as time varies. Unlike the usual projection of trajectories in the Lorenz system, for example, no topology for the dynamics is evident in this figure. Indeed, for this value of $L$ it is not even obvious that the dynamics is chaotic.
@@ III.1 Proper orthogonal decomposition
The most common dimensionality-reduction techniques are linear methods based on eigen-analysis such as dynamic mode decomposition (DMD) (Schmid, 2022) and principal component analysis (PCA) (Pearson, 1901; Hotelling, 1936), better known as proper orthogonal decomposition (POD) in the fluid dynamics community. POD consists of determining the dominant (uncorrelated) modes that capture most of the system’s variance, or ‘kinetic energy’. Given a timeseries $\{\bm{u}_{i}\}_{i=1}^{r}$ with $r$ time-steps, where $\bm{u}_{i}\in\mathbb{R}^{N_{x}}$ we can calculate the POD modes by stacking the zero-mean timeseries $\bm{\tilde{u}}_{i}=\bm{u}_{i}-\bm{u_{\mathrm{mean}}}$ in a matrix $\bm{\tilde{U}}\in\mathbb{R}^{r\times N_{x}}$ (where the rows are $r$ time-steps), and by considering its covariance matrix. The unbiased estimator $\bm{C}\in\mathbb{R}^{N_{x}\times N_{x}}$ for the covariance matrix is given by
@@ III.2 Autoencoder
Since we define the latent dimension to be $N_{h}=3$, this gives us the desired 3-dimensional representation of the KSE, in which we would like to test the methods described in section IV.1. Taking datapoints from a long physical timeseries and applying the encoder, we obtain a 3D latent timeseries. Figure 6 shows a 3D plot of the image of the attractor in latent space, illustrating its band-like shape, and 2D projections.
Figure 6: 3D plot (top) and 2D projections (bottom) of the chaotic attractor in the autoencoder’s latent space. The latent coordinates $(h_{1},h_{2},h_{3})$ are obtained by applying the encoder $\mathcal{E}$ to a data point. The grey points are from a timeseries solving the PDE. The blue line is a periodic orbit with period $T\approx 51.62$, plotted to emphasise the shape of the attractor.
@@ III.3.1 POD return map
In the POD latent coordinates $(\xi_{1},\xi_{2},\xi_{3})$, we define a Poincaré section $\mathcal{P}_{\mathrm{POD}}$ given by the plane $\xi_{1}=1.3$ and flow direction $\partial_{t}\xi_{1}<0$. The value $\xi_{1}=1.3$ is chosen so that the flow is approximately orthogonal to the Poincaré section. This defines a unique crossing section along the band (see figure 7), and gives us a sequence $\{\xi_{2}^{(i)}\}$ of the $\xi_{2}$ coordinate of the crossings through $\mathcal{P}_{\mathrm{POD}}$. The return map $f_{1}$ is defined by $f_{1}:\xi_{2}^{(i)}\mapsto\xi_{2}^{(i+1)}$. A plot of such subsequent crossings is shown in the top left panel of figure 8. The points approximately fall on a smooth unimodal curve, which we fit with a polynomial of order three. This simplified first return map appears to only have one intersection with the identity $\xi_{2}^{(i)}=\xi_{2}^{(i+1)}$, suggesting that there is only one periodic orbit with $p=1$, where $p$ is the number of intersections of an orbit with $\mathcal{P}_{\mathrm{POD}}$. By looking at further iterations/compositions of the return map (see other panels of figure 8), we also obtain an indication that there are 1, 0 and 1 periodic orbits with $p=2,3,4$ respectively (accounting for crossings that correspond to periodic orbits with smaller $p$ e.g. the orbit with $p=1$ will also account for one intersection in all subsequent graphs). Finally, we split the band domain $I=[-3.229,-2.952]$ at the minimum value of the first return map, giving approximately $I_{1}=[-3.229,-3.055]$ and $I_{2}=[-3.055,-2.952]$. From this, we obtain an indication of the qualitative dynamics: points starting in $I_{1}$ are mapped to the full interval $I$, while points starting in $I_{2}$ are mapped to a subset of $I_{1}$. This is shown in figure 9.
Figure 7: $(\xi_{1},\xi_{2})$ projection of the chaotic attractor in the latent space defined by the POD coordinates. The grey points are from a physical timeseries that is put through the encoder. The magenta line illustrates the Poincaré section $\xi_{1}=1.3$ and $\partial_{t}\xi_{1}<0$, denoted $\mathcal{P}_{\mathrm{POD}}$.
@@
Figure 9: First return map for the Poincaré section $\mathcal{P}_{\mathrm{POD}}$. The blue dots show data from a chaotic timeseries, whereas the green dot, red line and black line show the periodic orbits $1_{1}$, $2_{1}$ and $4_{1}$ respectively (see table 1). Below the figure, the intervals represent the first and second level within the state space partition derived from the simplified return map shown in figure 11. These intervals are used to construct symbolic sequences for the periodic orbits.